In each reaction box, place the best reagent or re
Answer #1
Concepts and reason

• One or more substances reacting together to form a product is called a chemical reaction. Most chemical reactions take place with the help of reagents.

• Reagents are the substance used to transform one form of substance to another.

• The reaction undergoes a series of steps like deprotonation of an alkyne, substitution, and hydrogenation reactions to attain a final product.

Fundamentals

The reagents play a vital role in all the reactions. Each reagent plays a different role in the chemical reaction.

Deprotonation of alkynes: An alkyne undergoes deprotonation with the help of the reagent sodium amide (NaNH2)\left( {{\rm{NaN}}{{\rm{H}}_{\rm{2}}}} \right) . Acetylide ion forms when sodium amide reacts with an alkyne. Acetylide ion acts as a good nucleophile which undergoes a substitution reaction with alkyl halides and forms a new carbon chain.

Substitution: The nucleophilic substitution reaction (SN2)\left( {{{\rm{S}}_{\rm{N}}}{\rm{2}}} \right) takes place when the nucleophile undergoes reaction with alkyl halides.

Hydrogenation: The addition of H2{{\rm{H}}_{\rm{2}}} in the presence of palladium is used to reduce alkenes and alkynes to alkanes.

NaNH
HC=co Nat
H-C=—H
A acetylene

H-CEC: Na
→H-CEC-
в
1-bromo-3-methylbutane

NaNH),
Nå
Na
:
3-
=
H-CEC-

cyclohexanone

H, Pd/C

Ans:

A
acetylene
1.1) NaNH |
1.2) 1-bromo-3-methylbutane
H-CEC-
NaNH
cyclohexanone
2.2)
step 3 reagent
H, Pd/C